OpenStudy (anonymous):

Which is the absolute value function corresponding to the graph with these characteristics? Vertex is at (-2, k) V-shaped graph Opens in the downward direction Coefficient a is 4 Left arm passes through point (-1, -3)

OpenStudy (anonymous):

A)y = 4|x − 2| B)y = 4|x − 2| + 3 C)y = -4|x + 2| + 3 D)y = -4|x + 2| − 3

OpenStudy (anonymous):

@battlefront64